OpenStudy (curry):

prove limit does not exist with polar coordinates. lim (x,y) => (0,0) [(x^2y)/(x^2 + y^2)^(3/2)

OpenStudy (dumbcow):

x^2y = (r^2 cos^2)(r sin) = r^3 cos^2 sin (x^2 + y^2)^3/2 = (r^2)^3/2 = r^3 \[\lim_{(xy) \rightarrow (0,0)} \cos^2 \theta \sin \theta\] not quite sure where to go from here

jimthompson5910 (jim_thompson5910):

as (x,y) --> (0,0) this is the same as saying r --> 0

jimthompson5910 (jim_thompson5910):

I think the fact that theta being on the right side makes this limit not equal to some fixed number, which points to the limit not existing. Not sure though

OpenStudy (dumbcow):

could be .... tan theta = y/x = 0/0 so theta is indeterminate so limit is indeterminate